OpenStudy (anonymous):

I did a math square and i know that the first two rows are absolutely correct. For the second row i found matches for most of them but i know i'm doing something wrong. someone plz help I've been working on this thing for like the past 3 days.

OpenStudy (sirm3d):

what is a math square?

OpenStudy (sirm3d):

what's IN a math square?

OpenStudy (anonymous):

the math square is the picture i took.

OpenStudy (anonymous):

all the sides have to = each other

OpenStudy (anonymous):

the first two rows are correct row 4,6 and 3,1

OpenStudy (anonymous):

but the next two are a little off i think i made a small mistake but i cant figure it out

OpenStudy (sirm3d):

that's a lot of work to do.

OpenStudy (anonymous):

you dont have to do the work ill do it could you just please tell me what i messed up on?

OpenStudy (sirm3d):

top-left square: is this the correct equation? \[\left\{\begin{matrix}x-4y=-2 \\ -3x+8y=-1\end{matrix}\right.\]

OpenStudy (anonymous):

yes

OpenStudy (sirm3d):

and the solution is (-5, -3/4)?

OpenStudy (anonymous):

i got -5, 7/4 but i know i did the first two rows right its just the next two rows

OpenStudy (anonymous):

umm what do you get for 3x -y = 12 -x +8y = -4

OpenStudy (sirm3d):

uhm, my mistake. your answer is correct.

OpenStudy (sirm3d):

i'd like to write down the entries first, left-to-ritgh, top-to-bottom, N-E-S-W order. (1,1): 4,6 (5,7/4) NA [no answer], (-6,-7). am i correct?

OpenStudy (anonymous):

yes